For this reaction, 32.0 g bromine reacts with 12.5 g chlorine gas. bromine (g) + chlorine (g) bromine monochloride (g)

What is the maximum mass of bromine monochloride that can be formed?

What is the FORMULA for the limiting reagent?

What mass of the excess reagent remains after the reaction is complete?
Answer #1

Br2 + Cl2 -------------> 2BrCl

according to balanced reaction

70.9 g Cl2 reacts with 159.8 g Br2

12.5 g Cl2 reacts with 12.5 x 159.8 / 70.9 = 28.17 g Br2

but we have 32.0 g Br2. so Br2 is exess reagent

limiting reagent = Cl2.

ecess Br2 left = 32.0 - 28.17 = 3.82 g

now

70.9g Cl2 gives 2 x 115.36 g BrCl

12.5 g Cl2 gives 12.5 x 2 x 115.36 / 70.9 =40.68 g BrCl

mass of BrCl formed = 40.68 g